# Device-Independent Certification of Multipartite Distillable Entanglement

### v1

@@ Acknowledgements.
We thank Ashutosh Marwah for pointing us to [43, Appendix A]. We acknowledge funding from Air Force Office of Scientific Research Award No. FA8750-23-2-0031. We also acknowledge support from the School of Electrical and Computer Engineering, Cornell University. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes, notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.

### v2

@@ Acknowledgements.
We thank Ashutosh Marwah for pointing us to [44, Appendix A]. Part of this work was completed during the conference “Beyond IID in Information Theory,” held at the University of Illinois Urbana-Champaign from July 29 to August 2, 2024, and supported by NSF Grant No. 2409823. We thank Kaiyuan Ji, Dhrumil Patel, Vishal Singh, and Theshani Nuradha for helpful discussions. We acknowledge funding from Air Force Office of Scientific Research Award Nos. FA9550-20-1-0067 and FA8750-23-2-0031. We also acknowledge support from the School of Electrical and Computer Engineering, Cornell University. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes, notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.